OpenStudy (anonymous):

Which of the following expressions are equivalent? Justify your reasoning. A. 4√x3 B. 1/(x^-1) C. 10√x5•x4•x2 D. x^1/3* x^1/3* x^1/3

OpenStudy (anonymous):

a little challenge for you guys

OpenStudy (anonymous):

correct answer get medal and fan

OpenStudy (danjs):

Do you know the exponent properties?

OpenStudy (danjs):

B and D are the same

OpenStudy (danjs):

X

OpenStudy (anonymous):

can you justify why

OpenStudy (danjs):

\[\frac{ 1 }{ x^{-1} } = x^1 = x\] \[x^{\frac{ 1 }{ 3 }+\frac{ 1 }{ 3 }+\frac{ 1 }{ 3 }} = x^\frac{ 3 }{ 3} = x^1 = x\]

OpenStudy (danjs):

hah, earning medals

OpenStudy (anonymous):

yes you are lol

OpenStudy (danjs):

I wont again, since it is a challenge for someone

OpenStudy (anonymous):

lol iguess